Omni Telecom is trying to decide whether to increase its cash dividend immediately or use the funds to increase its future growth rate.

   

P0 =

D1

Keg

P0 = Price of the stock today
D1 = Dividend at the end of the first year
D1 = D0 × (1 + g)
D0 = Dividend today
Ke = Required rate of return
g = Constant growth rate in dividends


D0 is currently $2.90, Ke is 9 percent, and g is 6 percent.

Under Plan A, D0 would be immediately increased to $3.30 and Ke and g will remain unchanged.
Under Plan B, D0 will remain at $2.90 but g will go up to 7 percent and Ke will remain unchanged.

a. Compute P0 (price of the stock today) under Plan A. Note D1 will be equal to D0 × (1 + g) or $3.30 (1.06). Ke will equal 9 percent, and g will equal 6 percent. (Round your intermediate calculations and final answer to 2 decimal places.)
Stock price for plan A =

b. Compute P0 (price of the stock today) under Plan B. Note D1 will be equal to D0 × (1 + g) or $2.90 (1.07). Ke will be equal to 9 percent, and g will be equal to 7 percent. (Round your intermediate calculations and final answer to 2 decimal places.)
Stock price for plan B =

c. Which plan will produce the higher value?

Plan A
Plan B
Answer #1

a.

D1 = 3.30*(1+0.06) = 3.498

P0 = D1/(Ke-g) = 3.498/(0.09-0.06) = $116.60

b.

D1 = 2.90*(1+0.07) = 3.103

P0 = 3.103/(0.09-0.07) = $155.15

c.

Plan B has higher value.